OpenStudy (anonymous):

2 7 -- + -- 5x 10x Is it A) 11 --- 10x OR B) 11 -- 10x^2

OpenStudy (akashdeepdeb):

Find the LCM here 2/5x + 7/10x \[\frac{ 2 }{ 5x } + \frac{ 7 }{ 10x }\] NOW FIND LCM by multiplying first term by 2 \[\frac{ 2*2 }{ 5x * 2 } + \frac{ 7 }{ 10x }\] NOW JUST CALCULATE NORMALLY BECAUSE THE DENOMINATOR IS THE SAME!! :D \[\frac{ 4 }{ 10x } + \frac{ 7 }{ 10x }\] \[ = \frac{ 11 }{ 10x }\] Understood this? :)